Job separation rate: .04

Job finding rate: .10

Labor force: 100 million

In the steady state:

what is the unemployment rate?

How many individuals lose their job each month?

How many individuals find jobs each month?
Answer #1

In steady state, unemployment rate (U/L) = s/(s+f) = 0.04/(0.04 + 0.10) = 0.04/0.14 = 0.2857 = 28.57%

Labor force = 100 millions

No. Of unemployed = (100 * 0.2857) millions = 28.57 millions

No. Of employed = (labor force - no. Of unemployed) = 71.43 millions

Individuals lose their job each year = (71.43 * 0.04) millions = 2.8572 millions

Individuals find jobs each month = (28.57 * 0.10) millions = 2.857 millions